OpenStudy (anonymous):

suppose you wish to tile a floor that measures 18 ft by 21 ft. how many square tiles that measure 3ft on a side should you purchase for the job?

OpenStudy (dumbcow):

Total area of floor is 21*18 Area of tile is 3^2 = 9 Number of tiles needed is Total Area/area of tile _> N = (21*18)/9 = 21*2 = 42